OpenStudy (lilai3):

A jar holds 15 red pencils and 10 blue pencils. One pencil is chosen, not replaced, and another pencil is chosen. What is the probability of drawing 2 red pencils from the jar?

OpenStudy (anonymous):

15/25*14/24 14 red pencils left and 24 total on the second pick

OpenStudy (lilai3):

210/600?

OpenStudy (anonymous):

Yep, and you can simplify that fraction as well. 210 = 3*10*7 600=3*10*20, so bottom line, 7/20

OpenStudy (lilai3):

thank you, @m_charron2! <333 :]